# EarthMarker: Visual Prompt Learning for Region-level and Point-level Remote Sensing Imagery Comprehension

### v2

@@ V-A Implementation Details
The proposed EarthMarker adopts the cross-domain phased training strategy, and the parameters updated vary at different stages. In general, we train an off-the-shelf 13B language model Llama 2 and the visual encoder MoV is kept frozen during the training. In the first multi-domain image-text alignment stage, only the alignment projection layer is updated. Then, in the spatial perception tuning phase, only the attention layers of LLM are unfrozen. Furthermore, the trainable LoRA metrics are introduced in the last RS visual prompting tuning stage. We utilized AdamW optimizer[65] with weight decay = 0 and betas = (0.9, 0.95), the learning rate is set to 2e-5, and the total training stages are conducted on 8 NVIDIA A100 GPUs. For model evaluation, we select diverse multi-granularity visual tasks to assess the performance of EarthMarker. Image-level tasks include scene classification and image captioning, while region-level tasks contain referring object classification, region captioning, and relationship analyses.
Methods | Publication Year | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE-L | CIDEr(0-5) | SPICE |
Expert Models | |||||||||
CSMLF [70] | GRSL 2019 | 77.0 | 64.9 | 53.2 | 47.1 | 32.0 | 57.8 | 106.5 | 26.5 |
Qu, et. al [43] | CITS 2016 | 72.5 | 60.3 | 51.8 | 45.5 | 33.6 | 59.1 | 117.9 | 27.6 |
Attention(soft) [71] | TGRS 2017 | 73.1 | 60.9 | 52.5 | 46.2 | 33.9 | 59.9 | 113.6 | 28.5 |
Attention(hard) [71] | TGRS 2017 | 73.3 | 61.0 | 52.7 | 46.4 | 34.0 | 60.0 | 110.3 | 28.4 |
FC-Att [72] | TGRS 2019 | 73.6 | 61.5 | 53.2 | 46.9 | 33.8 | 60.0 | 123.1 | 28.3 |
SM-Att [72] | TGRS 2019 | 73.9 | 61.7 | 53.2 | 46.8 | 33.0 | 59.3 | 123.6 | 27.6 |
MLCA-Net [73] | TGRS 2022 | 74.5 | 62.4 | 54.1 | 47.8 | 33.7 | 60.1 | 126.4 | 28.5 |
Visual Prompting Model | |||||||||
@@ V-B Scene Classification
For scene classification tasks, we use the AID[74] and UCMerced[75] datasets for evaluation. AID is a large-scale aerial dataset collected from Google Earth, containing 30 categories. Following the setting of GeoChat, we use a 20 split of the AID dataset for testing. The UCMerced dataset consists of 21 categories for scene classification. Following the setting of GeoChat, the entire UCMerced dataset is adopted as a zero-shot test set.
We prompt the model with an image-level box to represent the entire image. The text instruction is “Please identify the object category of each marked region in the image”. We calculate the zero-shot accuracy on the AID and UCMerced dataset. EarthMarker significantly outperforms other VLMs, with an accuracy of 86.52 on UCMerced and 77.97 on AID, as presented in Tab. II. In comparison, LLaVa-1.5 and Sphinx, due to the lack of RS domain knowledge, are inferior to the RS MLLM GeoChat and our EarthMarker. Compared to GeoChat, our EarthMarker achieved an accuracy improvement of 5.94 on AID and 2.09 on UCMerced. Owing to the multi-domain image-text alignment training, EarthMarker is endowed with excellent holistic scene understanding ability on RS imagery.
Method | Publication Year | Formats | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDER | SPICE |
MLLM | ||||||||||
Qwen-VL-Chat [69] | Arxiv 2023 | Coor | 20.66 | 9.55 | 4.90 | 1.96 | 8.35 | 20.93 | 29.18 | 9.55 |
GeoChat [66] | CVPR 2024 | Coor | 20.86 | 9.63 | 5.43 | 3.25 | 12.94 | 26.98 | 30.92 | 24.97 |
Sphinx [27] | Arxiv 2023 | Coor | 43.32 | 33.58 | 27.58 | 22.81 | 21.70 | 47.81 | 235.07 | 38.12 |
EarthGPT [5] | TGRS 2024 | Coor | 49.73 | 39.24 | 32.50 | 26.79 | 24.09 | 47.87 | 232.79 | 38.29 |
Visual Prompting Model | ||||||||||
Sphinx-V [15] | Arxiv 2024 | Box | 45.19 | 34.83 | 28.59 | 23.64 | 24.34 | 50.54 | 235.09 | 43.53 |
Vip-LLava-7b [76] | CVPR 2024 | Box | 21.03 | 11.26 | 6.37 | 3.25 | 9.59 | 23.50 | 30.95 | 12.44 |
Vip-LLava-13b [76] | CVPR 2024 | Box | 21.68 | 11.17 | 6.23 | 3.38 | 9.18 | 23.22 | 28.58 | 10.68 |
GLaMM [76] | CVPR 2024 | Box | 23.30 | 14.18 | 8.78 | 5.15 | 11.24 | 30.36 | 64.08 | 17.99 |
EarthMarker(Ours) | Point | 57.45 | 49.23 | 43.88 | 39.46 | 32.26 | 61.51 | 400.76 | 61.03 | |
@@
Methods | Publication Year | Formats | SS | SIOU |
MLLMs | ||||
GeoChat [69] | CVPR 2024 | Coor | 79.59 | 68.80 |
Sphinx [27] | Arxiv 2023 | Coor | 93.72 | 89.37 |
EarthGPT [5] | TGRS 2024 | Coor | 94.64 | 90.16 |
Visual Prompting Models | ||||
Sphinx-V [15] | Arxiv 2024 | Box | 89.07 | 81.62 |
Vip-LLava-7b [76] | CVPR 2024 | Box | 72.56 | 55.94 |
Vip-LLava-13b [76] | CVPR 2024 | Box | 74.51 | 60.53 |
EarthMarker(Ours) | Point | 95.96 | 93.49 | |
EarthMarker(Ours) | Box | 98.37 | 97.24 |
@@ V-C Image Captioning
To evaluate the image captioning capabilities, we use the NWPU-Captions[41] dataset to assess and compare EarthMarker against other expert models in the supervised setting. Created by Northwestern Polytechnical University, the NWPU-Captions dataset includes 31,500 aerial images and 157,500 sentences for RS image description. Following the protocol of MLCA-Net, we employ BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and CIDErD as evaluation metrics. In the evaluation, we use the sentence “Please provide a brief caption of each marked region in the image.” as text instruction and a full-image box as the visual prompt. As shown in Tab. III, compared to other expert models, EarthMarker demonstrates improvements in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, and ROUGE-L by 7.4, 8.2, 8.8, 6.5, 3.5, and 9.9, respectively, and a 36.5 improvement in CIDErD. In summary, we dexterously annotate the entire image as a single region, and further using both natural and RS image training, EarthMarker achieves a high level of understanding of the overall image, outperforming expert models.
@@ V-D Referring Object Classification
The referring object classification task aims to identify the category within the referring region in the image. The metrics used to evaluate this task are two semantic relevance indicators—Semantic Similarity (SS) and Semantic Intersection over Union (S-IOU) to assess a model’s classification ability. Closed-set testing is adopted on the test sets of object-level dataset DIOR-RSVG [19]. The text instruction is “Please identify the category of the marked region in the image”, which along with the bounding boxes are fed into LLM to predict the category of regions. Due to the former MLLM (e.g., GeoChat, Sphinx, and EarthGPT) only accepting images and text as input, the region prompt for those MLLMs are coordinates information contained in the text instructions.
As the results shown in Tab. V, EarthMarker achieves 95.96 in SS and 93.49 in S-IoU using point-level visual prompts, and 98.37 in SS and 97.24 in S-IoU based on box-level visual prompts on the DIOR-RSVG dataset. Both sets of results significantly outperform the SOTA method. Furthermore, EarthMarker surpasses the previous SOTA model EarthGPT by 3.73 in SS and 7.08 in S-IoU on the DIOR-RSVG dataset, demonstrating its superior capability in fine-grained box-level classification and the effectiveness of our designed visual prompting method. Additionally, in Fig. 3, there are considerable differences in predictions of regional object categories by EarthMarker and other MLLMs, as well as visual prompting models. It is evident that when faced with complex geographical scenes and blurry tiny objects, EarthMarker’s predictions are significantly accurate.
@@ V-E Region Captioning
For the brief region captioning, the test set of DIOR-RSVG [19] is employed. Specifically, we adopt boxes as the visual prompt and a text prompt, such as “Please provide a brief caption of each marked region in the image,” to prompt EarthMarker to concisely describe the content of the specified region using a brief caption. Similar to the image captioning task, metrics like BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and SPICE are used to evaluate EarthMarker and other MLLMs, visual promoting models in region target understanding.
As displayed in Tab. IV, on the DIOR-RSVG test set, compared to other MLLMs such as Qwen-VL-Chat, GeoChat, SPhinx, EarthGPT, and visual prompting models such as Sphinx-V, ViP-LLava, and GLAMM, EarthMarker shows improvements of 7.72, 9.99, 11.38, 12.67, blue12.49, 7.92, 10.97, 165.67, and 17.50 in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, CIDER and SPICE, respectively. Furthermore, we visualize the instances of EarthMarker and other models on the region captioning task, as shown in Fig. 4. In complex RS scenarios with numerous targets and extensive geographic coverage, EarthMarker can accurately identify and describe various specified targets, which are challenging tasks for other models. For example, compared to the proposed EarthMarker, the current popular MLLMs like GeoChat and EarthGPT are still inferior in regional comprehension. This excellent performance is mainly attributed to the region-text training and the dual instructions design of visual prompts and language.
V-F Complex Visual Reasoning

### v3

@@ V-A2 Implementation Details
The proposed EarthMarker adopts the cross-domain training strategy, and the parameters updated vary at different stages. In general, we train an off-the-shelf 13B language model Llama 2 and the visual encoder MoV is kept frozen during the training. In the first multi-domain image-text alignment stage, only the alignment projection layer is updated. Then, in the spatial perception tuning phase, only the attention layers of LLM are unfrozen. Furthermore, the trainable LoRA metrics are introduced in the last RS visual prompting tuning stage. We utilize AdamW optimizer[89] with weight decay = 0 and betas = (0.9, 0.95), the learning rate is set to 2e-5, and the total training stages are conducted on 8 NVIDIA A100 GPUs (each equipped with 80 GB of memory) and the training time is around 316 hours.
For model evaluation, we select diverse multi-granularity visual tasks to assess the performance of EarthMarker. Image-level tasks include scene classification and image captioning, while region-level tasks contain referring object classification, region captioning, and relationship analyses. In addition, we evaluate EarthMarker under both zero-shot and supervised settings to compare its performance with other existing models. To fully validate the superiority of EarthMarker, we select a range of models, including natural scene MLLMs, RS-specific MLLMs, and expert models. Note that the newly constructed RSVP contains only training samples for developing visual prompting MLLMs and does not include a test set. For the test sets, we use new datasets not involved in RSVP or test subsets from raw datasets, depending on the specific task, to ensure a comprehensive evaluation. Specifically, in the scene classification task, we use a zero-shot setting with new datasets such as AID [90] and UCMerced [91]. For other tasks, we use a supervised setting and evaluate the model performance on the test sets of DIOR-RSVG, etc. More descriptions of model selection and settings are detailed in Sections V-B to V-E.
Models | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE-L | CIDEr(0-5) | SPICE |
Expert Models | ||||||||
CSMLF [95] | 77.0 | 64.9 | 53.2 | 47.1 | 32.0 | 57.8 | 106.5 | 26.5 |
Qu, et. al [61] | 72.5 | 60.3 | 51.8 | 45.5 | 33.6 | 59.1 | 117.9 | 27.6 |
Attention (soft) [96] | 73.1 | 60.9 | 52.5 | 46.2 | 33.9 | 59.9 | 113.6 | 28.5 |
Attention (hard) [96] | 73.3 | 61.0 | 52.7 | 46.4 | 34.0 | 60.0 | 110.3 | 28.4 |
FC-Att [97] | 73.6 | 61.5 | 53.2 | 46.9 | 33.8 | 60.0 | 123.1 | 28.3 |
SM-Att [97] | 73.9 | 61.7 | 53.2 | 46.8 | 33.0 | 59.3 | 123.6 | 27.6 |
MLCA-Net [98] | 74.5 | 62.4 | 54.1 | 47.8 | 33.7 | 60.1 | 126.4 | 28.5 |
Visual Prompting Model | ||||||||
@@ V-B Scene Classification
For scene classification tasks[99], we use the AID[90] and UCMerced[91] datasets for evaluation. AID is a large-scale aerial dataset collected from Google Earth, containing 30 categories. Following the setting of GeoChat[6], we use a 20 split of the AID dataset for testing. Moreover, the entire UCMerced dataset is adopted as a zero-shot test set, which consists of 21 categories for scene classification.
We prompt the model with an image-level box to represent the entire image. The text instruction is “Please identify the object category of each marked region in the image”. We calculate the zero-shot accuracy on the AID and UCMerced dataset. EarthMarker significantly outperforms other MLLMs, with an accuracy of 86.52 on UCMerced and 77.97 on AID, as presented in Tab. II. In comparison, LLaVa-1.5 and Sphinx, due to the lack of RS domain knowledge, are inferior to the RS MLLM GeoChat and our EarthMarker. Compared to GeoChat, our EarthMarker achieves an accuracy improvement of 5.94 on AID and 2.09 on UCMerced. Owing to the multi-domain image-text alignment training, EarthMarker is endowed with excellent holistic scene understanding ability on RS imagery.
Models | Formats | SS | SIOU |
MLLMs | |||
GeoChat [6] | Coor | 79.59 | 68.80 |
Sphinx [31] | Coor | 93.72 | 89.37 |
EarthGPT [18] | Coor | 94.64 | 90.16 |
Visual Prompting Models | |||
Sphinx-V [16] | Box | 89.07 | 81.62 |
Vip-LLava-7b [100] | Box | 72.56 | 55.94 |
Vip-LLava-13b [100] | Box | 74.51 | 60.53 |
EarthMarker (Ours) | Point | 95.96 | 93.49 |
EarthMarker (Ours) | Box | 98.37 | 97.24 |
Models | Formats | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDEr | SPICE |
MLLMs | |||||||||
Qwen-VL-Chat [6] | Coor | 20.66 | 9.55 | 4.90 | 1.96 | 8.35 | 20.93 | 29.18 | 9.55 |
GeoChat [92] | Coor | 20.86 | 9.63 | 5.43 | 3.25 | 12.94 | 26.98 | 30.92 | 24.97 |
Sphinx [31] | Coor | 43.32 | 33.58 | 27.58 | 22.81 | 21.70 | 47.81 | 235.07 | 38.12 |
EarthGPT [18] | Coor | 49.73 | 39.24 | 32.50 | 26.79 | 24.09 | 47.87 | 232.79 | 38.29 |
Visual Prompting Models | |||||||||
Sphinx-V [16] | Box | 45.19 | 34.83 | 28.59 | 23.64 | 24.34 | 50.54 | 235.09 | 43.53 |
Vip-LLava-7b [100] | Box | 21.03 | 11.26 | 6.37 | 3.25 | 9.59 | 23.50 | 30.95 | 12.44 |
Vip-LLava-13b [100] | Box | 21.68 | 11.17 | 6.23 | 3.38 | 9.18 | 23.22 | 28.58 | 10.68 |
GLaMM [100] | Box | 23.30 | 14.18 | 8.78 | 5.15 | 11.24 | 30.36 | 64.08 | 17.99 |
EarthMarker(Ours) | Point | 57.45 | 49.23 | 43.88 | 39.46 | 32.26 | 61.51 | 400.76 | 61.03 |
@@
CNN Model | ViT Model | Classification | Region Captioning | Referring Object Classification | |
CLIP ConvNeXt-L | CLIP RN5016 | DINOv2 ViT-L/14 | AID (Top-1 Acc) | DIOR-RSVG (CIDEr) | DIOR-RSVG (SIOU) |
✓ | 73.83 | 368.21 | 94.27 | ||
✓ | 74.09 | 368.10 | 94.52 | ||
✓ | 73.21 | 369.84 | 95.93 | ||
✓ | ✓ | 75.39 | 372.41 | 96.81 | |
✓ | ✓ | 77.97 | 379.25 | 97.24 |
Training Data Composition | Classification | Region Captioning | Referring Object Classification | |||
Natural Caption | Natural Referring | RS Reconstructed | GPT4V Generated | AID (Top-1 Acc) | DIOR-RSVG (CIDEr) | DIOR-RSVG (SIOU) |
✓ | 75.24 | 352.82 | 96.46 | |||
✓ | ✓ | 76.97 | 356.25 | 97.01 | ||
✓ | ✓ | ✓ | 77.94 | 358.50 | 97.20 | |
✓ | ✓ | ✓ | ✓ | 77.97 | 379.25 | 97.24 |
@@ V-C Image Captioning
To evaluate the image captioning capabilities, we use the NWPU-Captions[59] dataset to assess and compare EarthMarker against other expert models in the supervised setting. Created by Northwestern Polytechnical University, the NWPU-Captions dataset includes 31,500 aerial images and 157,500 sentences for RS image description. Following the protocol of MLCA-Net [98, 18], we employ BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and CIDErD as evaluation metrics, which offers a comprehensive language quality evaluation. In the evaluation, we use the sentence “Please provide a brief caption of each marked region in the image.” as text instruction and a full-image box as the visual prompt. As shown in Tab. III, compared to other expert models, EarthMarker demonstrates improvements in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, and ROUGE-L by 7.4, 8.2, 8.8, 6.5, 3.5, and 9.9, respectively, and a 36.5 improvement in CIDErD. In summary, we dexterously annotate the entire image as a single region, and further using both natural and RS image training, EarthMarker achieves a high level of understanding of the overall image, outperforming expert models.
@@ V-D Referring Object Classification
The referring object classification task aims to identify the category within the referring region in the image. Following [16, 14], the metrics used to evaluate this task are two semantic relevance indicators—Semantic Similarity (SS) and Semantic Intersection over Union (S-IOU) to assess a model’s object classification ability. Closed-set testing is adopted on the test sets of object-level dataset DIOR-RSVG [23]. The text instruction is “Please identify the category of the marked region in the image”, which along with the bounding boxes are fed into LLM to predict the category of regions. Due to the previous MLLM (e.g., GeoChat, Sphinx, and EarthGPT) only accepting images and text as input, the region prompts for those MLLMs are coordinates information contained in the text instructions.
As the results shown in Tab. IV, EarthMarker achieves 95.96 in SS and 93.49 in S-IoU using point-level visual prompts, and 98.37 in SS and 97.24 in S-IoU based on box-level visual prompts on the DIOR-RSVG dataset. Both sets of results significantly outperform the SOTA method. Furthermore, EarthMarker surpasses the previous SOTA model EarthGPT by 3.73 in SS and 7.08 in S-IoU on the DIOR-RSVG dataset, demonstrating its superior capability in fine-grained box-level classification and the effectiveness of our designed visual prompting method.
@@ V-E Region Captioning
For the brief region captioning, the test set of DIOR-RSVG [23] is employed. Specifically, we adopt boxes as the visual prompt and a text prompt, such as “Please provide a brief caption of each marked region in the image,” to prompt EarthMarker to concisely describe the content of the specified region using a brief caption. Similar to the image captioning task, metrics like BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and SPICE are used to evaluate EarthMarker and other MLLMs, visual promoting models in region target understanding. As displayed in Tab. V, on the DIOR-RSVG test set, compared to other MLLMs such as Qwen-VL-Chat, GeoChat, SPhinx, EarthGPT, and visual prompting models such as Sphinx-V, ViP-LLava, and GLAMM, EarthMarker shows improvements of 7.72, 9.99, 11.38, 12.67, 12.49, 7.92, 10.97, 165.67, and 17.50 in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, CIDER and SPICE, respectively.
V-F Ablation Analysis
Visual Encoders. To investigate the impact of our proposed mixture of visual experts (MOV) strategy on performance, we conduct an ablation experiment, and the results are presented in Tab. VI. It is clear that the designed MoV achieves the best performance across downstream tasks, compared with solely CNN or ViT[101] visual backbones. Specifically, when compared to using only CNN or ViT visual backbones, the proposed MoV (CLIP ConvNeXt-L and DINOv2 ViT-L/14) enhances performance in scene classification, region captioning, and referring object classification tasks by 3.88, 9.41, and 1.31, respectively. Furthermore, when compared to the combination of CLIP RN5016 and DINOv2 ViT-L/14, our designed MoV (CLIP ConvNeXt-L and DINOv2 ViT-L/14) improve performance in scene classification, region captioning, and referring object classification by 2.58, 6.84, and 0.43, respectively. This ablation experiment demonstrates the performance improvement of the MOV module for the proposed EarthMarker.
Training Data. We conduct ablation experiments on the training data to investigate the impact of different domains and GPT4V-assisted generated data on the model’s final performance. As shown in Tab. VII, leveraging natural scene captions and referring data significantly enhances the model’s performance across various visual tasks, including scene classification, region captioning, and referring object classification. Moreover, incorporating GPT4V-assisted generated data results in the best model performance, particularly showing notable improvements in the region captioning task. Therefore, multi-domain datasets and the long caption data synthesized by GPT4V enable EarthMarker to deliver accurate detailed interpretation of objects in images.
V-G Computation Analysis and Comparisons
In this part, we compare the GPU memory usage (MiB) and inference time (seconds) of EarthMarker with other MLLMs and visual prompting models in Tab. VIII. Specifically, we compare these two aspects of different models on the region captioning task using an NVIDIA RTX A6000. Note that for MLLMs and visual prompting models, GPU memory consumption is primarily driven by storing model weights. As shown in Tab. VIII, EarthMarker’s GPU memory consumption is larger than that of Sphinx [16], GeoChat[6], EarthGPT [18], Qwen-VL-Chat [92], and Vip-LLava-13b [100], but lower than that of Sphinx-V [16]. Regarding inference time, similarly, EarthMarker has a longer time consumption than all models except for Sphinx-V, due to its larger model parameters. Fortunately, the difference is mostly within 1 second, so the gap is not significant. In our future work, we will focus on optimizing EarthMarker’s inference time, aiming to strike a balance between inference speed and interpretative accuracy.
Models | Inference Time(s) | GPU Memory(MiB) |
MLLMs | ||
Sphinx | 1.08 | 41853 |
EarthGPT | 0.79 | 41766 |
GeoChat | 0.74 | 15544 |
Qwen-VL-Chat | 0.35 | 38145 |
Visual Prompting Models | ||
Sphinx-V | 1.95 | 41794 |
Vip-LLava-13B | 0.84 | 27748 |
EarthMarker(Ours) | 1.63 | 41038 |
VI VISUALIZATION
VI-A Comparison Between EarthMarker and Other Models
VI-A1 Referring Object Classification
In Fig. 3, there are considerable differences in predictions of regional object categories by EarthMarker and other MLLMs, as well as visual prompting models. In the top traffic scene image, EarthMarker correctly identifies the “Overpass” and the “Train station”, whereas other RS-specific MLLMs like GeoChat[6] and EarthGPT[18] misclassify the “Overpass” as a “Bridge” and the “Train Station” as “Storage Tanks”. Other visual prompting models, like Sphinx-V[16], mistakenly identify Region 2 as a “Expressway-service-area”. In the port scene below, only EarthMarker completely interprets the small targets (i.e., car and ship). It is evident that when faced with complex geographical scenes and blurry tiny objects, EarthMarker’s predictions are significantly accurate.
VI-A2 Region Captioning
We visualize the instances of EarthMarker and other models on the region captioning task, as shown in Fig. 4. In complex RS scenarios with numerous, small targets, and extensive geographic coverage, EarthMarker can accurately identify and describe various specified targets (e.g., overpass, tennis court, airport, and stadium). Notably, in the intricate fourth image, EarthMarker can identify small targets, such as expressway toll stations, that are difficult for the human eye to discern. EarthMarker achieves this by integrating information from the surrounding environment, demonstrating an impressive ability to handle both the overall context and detailed local targets. However, interpreting those cases is challenging for other existing models. For example, the current popular MLLMs in the RS domain like GeoChat[6] and EarthGPT[18] are still inferior in regional comprehension of RS imagery compared to our proposed EarthMarker. Natural scene MLLMs Sphinx[31], visual prompting models Vip-LLava[100] and GLaMM[102] struggle to understand complex RS scenarios and interpret those small objects.
Thanks to the design of EarthMarker’s joint prompting design that integrates both visual and language prompts, and the construction of region-level and point-level visual prompting datasets, EarthMarker achieves exceptional performance in comprehending complicated RS scenarios and analyzing multiple small objects.
VI-B Complex Visual Reasoning
VI-B1 Low-resolution Scenarios Understanding
The aforementioned cases highlight EarthMarker’s strong potential for visual understanding of high-resolution imagery. In this section, we visualize EarthMarker’s performance across diverse scenes and varying spatial resolutions. As illustrated in Fig. 5, we present some image cases featuring diverse low resolutions. These images exhibit challenges such as scene blurriness, color distortion, and noise interference, which make manual interpretation relatively difficult. Notably, EarthMarker can still accurately classify objects in diverse scenarios (e.g., foggy harbor environments, urban scenes, and blurred industrial areas), based on the provided region/point-level visual prompts. It can also deliver brief or detailed descriptions of the objects or scenes and analyze the relationships among multiple specified objects. In summary, EarthMarker consistently demonstrates its exceptional capabilities when adapting to new test environments and interpreting scenes with varying spatial resolutions and diverse data sources.
VI-B2 Relationship Analyses Tasks